Payback Period and Accounting Rate of Return: Equal Annual Operating Cash Flows with Disinvestment
Minn is considering an investment proposal with the following cash flows:

Initial investment-depreciable assets $100,000
Net cash inflows from operations (per year for 10 years) 25,000
Disinvestment-depreciable assets 19,000

For parts b. and c., round answers to three decimal places, if applicable.

a. Determine the payback period.
Answer: 4 years

b. Determine the accounting rate of return on initial investment.

c. Determine the accounting rate of return on average investment.
Answer #1

B)Depreciation expense =[cost - residual value ]/useful life

                        = [100000 - 19000]/10

                        = 81000/10

                      = 8100

Annual Net Income= Net cash inflow - Deprecation expense

                         = 25000 -8100

                         = 16900

Accounting rare of return on Initial investment =Annual net income /Initial cost

                           = 16900 / 100000

                           = .169 or 16.90%

c)Accounting rate of return on average investment = Average net income over 10years /Average investment

                                          = 16900/59500

                                          = .28403 or 28.403%

working:

Average investment =[Beginning book value +ending book value]/2

                   = [100000+19000]/2

                   = 119000/2

                  = 59500